OpenStudy (anonymous):

The sum of Tony's age and his older brother's age is 29. The difference between their ages is 5 years. How old is Tony?

OpenStudy (anonymous):

@mathstudent55 help please

OpenStudy (akashdeepdeb):

Let Tony's age = x Older brother's age = y The sum of Tony's age and his older brother's age is 29 Means, x + y = 29 The difference between their ages is 5 years. Means, y - x = 5 [Because, y > x (Older brother is obviously "older" than Tony)] So, we have 2 equations now: y + x = 29 y - x = 5 Add these 2 equations. y + x + y - x = 34 2y = 34 y = 17 Hence, the older brother (y) = 17 years. Because y - x = 5 17 - x = 5 x = 12 Therefore, y = 17, x = 12. Hence, Tony is 12 years old. His older brother is 17 years old.